Solve by elimination: 4x+3y=3
X+3y=1

Answer by stanbon(75887)

Solve by elimination:
4x+3y=3
X+3y=1
---
Subtract and solve for "x"::
3x = 2
x = 2/3
-------------
Cheers,
Stan H.
-----------